OpenStudy (hellokitty17):

MEDALS HERE FOR RIGHT ANSWER!!! The cost to produce a product is modeled by the function f(x) = 5x2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. 5(x − 7)2 + 13; The minimum cost to produce the product is $13. 5(x − 7)2 + 13; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $258.

OpenStudy (hellokitty17):

@jcoury

oregonduck (oregonduck):

5x^2-70x+258 5x^2-70x = 5(x^2-14x) = 5(x^2-14x+49 -49) = 5( x-7)^2 - 245 5x^2-70x+258 = 5(x-7)^2 -245+258 5(x-7)^2+13 The vertex is (7,13) and the parabola slopes upward because the coefficient of x^2 is positive. Therefore, (7,13) is a minimum point 5(x − 7)2 + 13; The minimum cost to produce the product is $13.

oregonduck (oregonduck):

so a

OpenStudy (anonymous):

http://openstudy.com/study#/updates/5240a0c3e4b090015417d6b6 There you go

OpenStudy (anonymous):

nice copy paste skills :P

OpenStudy (hellokitty17):

Thanks boys!!! Medal eachother again

oregonduck (oregonduck):

lol thxs :)

OpenStudy (anonymous):

any more questions?

oregonduck (oregonduck):

1 more medal m8 for a title

OpenStudy (hellokitty17):

yea

OpenStudy (anonymous):

its says im offline XD

oregonduck (oregonduck):

yes

OpenStudy (anonymous):

my computer has serious issues

oregonduck (oregonduck):

lol

OpenStudy (anonymous):

its says your sister is offline too

oregonduck (oregonduck):

we have more questions so you can uncok your title

oregonduck (oregonduck):

yeah hers has some problems too

OpenStudy (anonymous):

don't forget to tag

oregonduck (oregonduck):

ok we won't

oregonduck (oregonduck):

we have 7 more questions

OpenStudy (anonymous):

k

OpenStudy (anonymous):

after that i have 4 questions i want you guys to answer

oregonduck (oregonduck):

ok

oregonduck (oregonduck):

try searching them first that is what i do lol

OpenStudy (anonymous):

i can't i have to take a picture

OpenStudy (anonymous):

that's what i do too

oregonduck (oregonduck):

someone i do not know just taged me that is how famous i am